As you can tell I am not to good with story problems and find this one confusing, can you help me figure it out.
The width of a rectangle is 1 ft less than the length. The area is 6ft^2. Find the length and the width.
The width is _____.
The length is _____.
Thank you so much. :o)

Answer by checkley77(12844)

W=L-1
A=LW
6=L(L-1)
6=L^2-L
L^2-L-6=0
(L-3)(L+2)=0
L-3=0
L=3 ANS. FOR THE LENGTH.
W=3-1=2 ANS. FOR THE LENGTH.
PROOF:
6=2*3
6=6